OpenStudy (highschoolmom2010):

A baseball team played 154 regular season games. The ratio of the number of games they won to the number of games they lost was 5/2. How many games did they win? How many games did they lose?

OpenStudy (jdoe0001):

well, it's saying they played a total of 154 games and they won 5/2 of that

ganeshie8 (ganeshie8):

neither

OpenStudy (highschoolmom2010):

win to loose was 5/2 sorry computer acting full rere mode

OpenStudy (jdoe0001):

hmm, I read it, of every 5 won, they lost 2 :/

ganeshie8 (ganeshie8):

so the ratio is \(\frac{5x}{2x}\) \(5x\) wins \(2x\) lost

ganeshie8 (ganeshie8):

\(5x+2x = 154\) solve x

OpenStudy (highschoolmom2010):

7/7x=154/7 x=22

ganeshie8 (ganeshie8):

correct ! so how many wins ?

OpenStudy (highschoolmom2010):

154-22= 132 :)

ganeshie8 (ganeshie8):

careful

OpenStudy (highschoolmom2010):

what did i do wrong?

OpenStudy (highschoolmom2010):